### v1

Abstract
We carried out long-term monitoring of the LIGO/Virgo/KAGRA binary black hole (BBH) merger candidate S230922g in search of electromagnetic emission from the interaction of the merger remnant with an embedding active galactic nuclei (AGN) accretion disk. Using a dataset primarily composed of wide-field imaging from the Dark Energy Camera (DECam) and supplemented by additional photometric and spectroscopic resources, we searched of the sky area probability for transient phenomena, and discovered 6 counterpart candidates. One especially promising candidate - AT 2023aagj - exhibited temporally varying asymmetric components in spectral broad line regions, a feature potentially indicative of an off-center event such as a BBH merger. This represents the first live search and multiwavelength, photometric, and spectroscopic monitoring of a GW BBH optical counterpart candidate in the disk of an AGN.
1 Introduction
As multimessenger astronomy continues to develop, the branch of this field dedicated to studying compact binary coalescences (CBCs) stands as an important forerunner as the first to include direct observations of gravitational waves (GWs) (Abbott et al., 2017a). These events manifest phenomena at scales presently unobtainable in man-made environments, and hence are important natural laboratories in which to study extreme physics such as heavy element nucleosynthesis (Symbalisty & Schramm, 1982) and measure key physical quantities such as the expansion rate of the universe (Schutz, 1986; Holz & Hughes, 2005).
This potential was realized following the first successful multimessenger detection of the binary neutron star merger GW170817/GRB 170817A (Abbott et al., 2017b), wherein a global electromagnetic astronomy campaign succeeded in locating a kilonova (KN) (Coulter et al., 2017; Arcavi et al., 2017; Andreoni et al., 2017; Drout et al., 2017; Nicholl et al., 2017; Evans et al., 2017; Lipunov et al., 2017; Pian et al., 2017; Smartt et al., 2017; Utsumi et al., 2017; Valenti et al., 2017; Kasliwal et al., 2017; Troja et al., 2017; Margutti et al., 2017; Shappee et al., 2017; Soares-Santos et al., 2017; Chornock et al., 2017; Tanvir et al., 2017) incident with the initially identified GW and gamma ray burst (GRB) events (Abbott et al., 2017a; Savchenko et al., 2017; Goldstein et al., 2017). This single event by itself was enough to result in hundreds of analyses, including constraints on neutron star (NS) physics (e.g. Abbott et al., 2018) and the Hubble constant (Abbott et al., 2017c). In the case of , it is predicted that binary neutron star (BNS) events with electromagnetic (EM) counterparts are required to make a few% measurement of the parameter (Chen et al., 2018), although joint multimessenger constraints on the binary viewing angle can further improve precision by a factor of a few (e.g. Hotokezaka et al. 2019; Palmese et al. 2024).
While BNS mergers are the only class of GW event with a confirmed EM counterpart, it is also predicted that binary black hole (BBH) mergers can produce counterparts in certain circumstances. The majority of proposed counterpart mechanisms involve the interaction of the binary or merger remnant with a coexistent medium, usually the gaseous accretion disk of an active galactic nucleus (AGN), whether through accretion (Bartos et al., 2017b), ram-pressure stripping of gas about the remnant or jetted Bondi accretion (McKernan et al., 2019), or breakout emission from a post-merger jet (Tagawa et al., 2023, 2024). A cartoon of the AGN-associated EM counterpart mechanism is show in Figure 1.
Counterpart candidates for BBH mergers from previous LIGO/Virgo/KAGRA (LVK) observing runs have been proposed (Graham et al., 2020, 2023), although confirmation of a counterpart has remained challenging. There are several reasons for this difficulty. First, even if we know that a given merger happens in an AGN, in principle there is only a chance that we could detect an EM counterpart: our view of around half of AGN (the Type 2s) is obscured, and an unfortunate kick-direction out the opposite side of a Type 1 AGN could still obscure any emission (Graham et al., 2023). Second, even if a flare emerges from an AGN, it must be discerned in the presence of other AGN variability. This makes our task of searching for counterparts more challenging in brighter AGN, and our search is biased against less luminous counterparts. Third, possible flare parameters are weakly constrained, both because AGN disk properties are uncertain to orders of magnitude, and because the properties of the post-merger flare depend on these uncertain parameters.
One possible means to confirm a counterpart is through optical spectroscopy: because the BBH merger occurs off-center in the AGN disk, the resulting flare could unevenly illuminate the broad-line region (BLR) of the AGN, resulting in asymmetric broad-line features in the optical spectrum (McKernan et al., 2019). Synchronized evolution of a transient with such spectral features can be a key piece of evidence in favor of the classification of a transient as a EM counterpart to a BBH merger. At the very least, an evolving asymmetric line profile suggests an off-center flaring event in the AGN, ruling out most sources of AGN variability near the SMBH. Flare energetics and lightcurve profiles can allow us to rule out e.g. embedded supernovae, leaving few candidates for a sufficiently energetic off-center AGN flare, including a BBH merger.
Certainly, the search for EM counterparts to BBH mergers has unique challenges versus the search for NS merger counterparts (intrinsic AGN variability, delay time uncertainties, etc.); however, as the historical rate of detection of BBH mergers has been greater than that for NS mergers, there are many more opportunities to search for a counterpart to the former kind of event. Because multimessenger observations of BBH mergers are also useful in making cosmological measurements (Palmese et al., 2021; Bom & Palmese, 2023; Alves et al., 2024), the pursuit of EM counterparts to BBH mergers has the potential to significantly contribute to the multimessenger observations required to make the first 2% measurement of the Hubble constant with standard sirens, which the level of precision needed to help us understand the Hubble tension (Chen et al., 2018).
Currently, GW follow-up of distant events is best enabled with technology capable of meeting the colloquial requirements of “wide, fast, deep”, as to ensure the rapid and thorough coverage of event volumes. The Dark Energy Camera (DECam) (Flaugher et al., 2015) on the 4m Victor M. Blanco Telescope st the Cerro Tololo Inter-American Observatory (CTIO) is one of the premiere instruments that can address these requirements. DECam already has a respectable history in GW follow-up, having been used for this purpose from the first GW event (Soares-Santos et al., 2016; Annis et al., 2016) and many events since then (Goldstein et al., 2019; Andreoni et al., 2019; Herner et al., 2020; Andreoni et al., 2020; Morgan et al., 2020; Garcia et al., 2020), and importantly being one of the first instruments to detect the counterpart KN to GW170817 (Soares-Santos et al., 2017). The new survey program Gravitational Wave MultiMessenger Astronomy DECam Survey (GW-MMADS) (PIs: Andreoni & Palmese) is designed to find EM counterparts to BNS, NS BH, and BBH mergers via rapid DECam follow-up of GW events during the fourth gravitational wave observing run (O4).
In this work, we present results from our follow up of the GW event S230922g. S230922g is a LVK BBH merger candidate detected at 2023-09-22 02:03:44.886 UTC (Ligo Scientific Collaboration et al., 2023a, b). The 90% credible region of 324 deg2 is one of the smallest areas of O4 a (the first part of O4, completed in January 2024); this, combined with the localization being visible from Chile in the months following the trigger, made an effective follow-up with DECam possible in pursuit of potential AGN-linked counterparts like those mentioned previously. In this paper we present our long-term monitoring of S230922g; the paper is organized as follows: §2 presents our program’s generic strategy for GW follow-up and the data collected, highlighting components especially relevant for S230922g; §3 describes the methods we used to distill our population of detected transients into a shortlist of counterpart candidates; §4 describes our most favored candidate, and to a lesser extent additional candidates of interest; in §5 we estimate parameters for our most favored counterpart candidate, finding it well within the confines of existing BBH counterpart theory; in §6 we summarize our findings and highlight considerations relevant for future efforts. Throughout this work we use a flat CDM model with km/s/Mpc and matter density of ,
2 Data
2.1 GW data
S230922g is an event of interest because it was detected with high significance (False Alarm Rate - FAR - 1 per years from gstlal; Messick et al. 2017; Sachdev et al. 2019) by both LIGO Livingston and LIGO Hanford, it is well localized compared to the O4a population, and it has 100% probability of being a BBH. It was also identified with high significance by the Burst CWB (Klimenko et al., 2016) search pipeline, potentially indicating a loud, short burst, as one may expect for a massive BBH spending a short fraction of the late inspiral phase in the LVK band. The luminosity distance of the event, marginalized over the entire sky, is Mpc. These quantities are reported with the skymap from the Bilby (Ashton et al., 2019) reduction (Ligo Scientific Collaboration et al., 2023c); we use this skymap for our work.
Higher mass () BBH mergers are more likely to have originated from the disks of AGNs (e.g. Gayathri et al. 2021), and so we also inform our decision to trigger BBH follow up by estimating the total mass of the binary. We follow a similar calculation to that in Graham et al. (2020), and consider that (Berry et al., 2015), as well as (Finn & Chernoff, 1993), where is the chirp mass of the binary. For this event, we assume a 1.4-1.4 binary neutron star merger detection horizon of 150 and 152 Mpc for Hanford and Livingston, respectively, an equal mass system, and the luminosity distance marginalized over the sky; we derive a total rest frame mass of for S230922g. Note that this estimate is highly uncertain with an error bar of at least a factor of 2.
2.2 DECam observations
Our team was notified of S230922g through a General Coordinates Network (GCN) listener333Adapted from https://github.com/scimma/slackbot. that sent a digest of the event to our Slack workspace.444Our team also uses the Fritz science data platform (Coughlin et al., 2023) for notifications of GW events via phone call, but we reserve this kind of notification for exceptionally time-sensitive events such as BNS and NS BH mergers. Our first response was to generate an initial observing strategy with gwemopt (Coughlin et al., 2018) to asses the probability coverage possible with a single night of DECam observations. gwemopt selects exposure pointings from a preset sky tiling. Because our image subtraction pipeline uses archival DECam exposures as templates, we implemented a custom tiling based on images calibrated by the NSF National Optical-Infrared Research Laboratory (NOIRLab) DECam Community Pipeline (CP) (Valdes et al., 2014) and made available through the NOIRLab Astro Data Archive555https://astroarchive.noirlab.edu/ (McManus & Olsen, 2021) to ensure our observing plans include pointings with usable templates. During each observing night, the CP calibrates the response and astrometry of each exposure shortly after it is taken which is then available to the image subtraction pipeline.
We determined that 70% of the GW localization area could be covered with 60 s and 80 s DECam exposures in - and -bands, respectively, with 4.7 hours of telescope time. Our DECam tiling is shown in Figure 2, overlaid on the LVK Bibly skymap (Ligo Scientific Collaboration et al., 2023b) for S230922g. When this plan was finalized, we communicated our pointings to the larger astronomical community though the GW Treasure Map666https://treasuremap.space/ (Wyatt et al., 2020), and we updated information appropriately as we executed our plan.
Leading BBH counterpart models predict delays of 1-3 weeks after the merger before the manifestation of any EM counterpart, due to the need for the signal to first emerge from the depths of the AGN disk. Even so, it is valuable to initiate EM follow up for these kinds of events within a few days of the GW trigger, as high-cadence archival data are generally not available in the localization area of a given GW event and a baseline of sources in the localization area can be useful when vetting candidates. Accordingly, we decided to trigger our Target of Opportunity (ToO) program for the night following the event (the evening of September 22, 2023) in order to establish a baseline of activity at the GW merger time and to facilitate the identification of novel phenomena in subsequent observing epochs.
On the first night we were only able to complete 1 hour of observations before inclement weather began, but we were able to observe our full plan in its entirety the following night (September 23). We publicly reported transients in the area through GCN within 24 hours of these observations (Cabrera et al., 2023). Further observations were conducted 11, 33, 41, and 70 days after the LVK trigger (October 3, October 25, November 2, and December 1). As we processed the data throughout the campaign, we took additional observations of the most interesting candidates, adding up to eight epochs altogether. We submitted all transient sources discovered through this campaign to the Fritz SkyPortal (van der Walt et al., 2019; Coughlin et al., 2023). After composing our final candidate shortlist (as detailed in §3), we reported the respective sources to the Transient Name Server (TNS).
2.3 Spectroscopic observations
BBH counterpart models have some degeneracy with those for stochastic AGN variability and other transient events like tidal distruption events (TDEs), but spectroscopic data has been predicted to serve as a key discriminator among these phenomena: specifically, the off-center location of a BBH merger in the accretion disk of the AGN as opposed to the central location of TDEs or accretion-based phenomena is expected to induce asymmetry in the broad lines of AGN spectra as the transient event “washes” over the BLR in an asymmetric manner (McKernan et al., 2019). The detection of a BLR asymmetry that evolves in concert with the light curve of a transient is a smoking gun that locates the transient in an off-center position in the accretion disk expected for BBH mergers, and the coincidence of such an event with a GW merger is a strong evidence standard astronomical techniques can provide linking BBH mergers and AGN flares.
In pursuit of such evidence, we triggered several spectroscopic resources for additional follow-up of transients whose light curves demonstrated proposed counterpart features (see §3.1). A total of five spectra were collected during our follow-up campaign: two with Gemini GMOS, and one each with Keck LRIS, SALT RSS, and P200 DBSP. We discuss the methodology for each spectrum in the context of the respective candidate in §4.2.
3 Methods
3.1 Automated vetting
Detection-based cuts | ||
Initial photometry detections | 25,392,140 | 1.00 |
Data quality masking | 14,053,259 | 0.553 |
Real/bogus score 0.7 | 1,501,800 | 0.059 |
Source-based cuts | ||
Initial transients | 233,313 | 1.00 |
Remove LS variable stars | 140,230 | 0.601 |
Remove MPC objects | 18,528 | 0.079 |
1 real/bogus score 0.9 | 17,515 | 0.075 |
2 detections with min | 3558 | 0.015 |
LS galaxy sep 1” | 2388 | 0.010 |
Additional vetting | 6 | 2.6e-5 |
We analyze our data with our difference photometry pipeline described in Hu et al. (2024). We perform image subtraction with Saccadic Fast Fourier Transform (SFFT) (Hu et al., 2022), a scalable image subtraction algorithm and GPU-enabled implementation of the same that produces difference images up to an order of magnitude faster than widely used tools with comparable accuracy. Aperture photometry is conducted on the resulting difference images using SExtractor (Bertin & Arnouts, 1996) with fluxes calibrated to the DESI Legacy Survey (LS) source catalog and corrected for extinction using the dustmaps package (Green, 2018) and the coefficients from Schlafly & Finkbeiner 2011. Our image differencing pipeline detected over 25 million transient features in our campaign data, which we distilled into a tractable list of astrophysical transients through a series of cuts. Table LABEL:tab:vetting summarizes the vetting steps we applied, which are described in this section.
We first remove any detections on difference images contaminated by bad pixels recorded in the DECam CP data quality mask products. Surviving features were then scored using a Convolutional Neural Network (CNN) trained on archival DECam data products from our pipeline to perform real/bogus classification. Possible CNN scores range from 0 to 1, with higher scores associated with more realistic astrophysical objects; at this step we kept any detections with scores greater than 0.7. Roughly 6% of the initial list of detections passed these steps to form our list of real astrophysical detections.
These detections were then cross-matched among themselves to identify features present in data from multiple epochs and filters; the resulting list of transients consisted of 233,313 objects. This list was subsequently crossmatched with the LS DR10.1 catalog (Dey et al., 2019) and the Minor Planet Center (MPC) service, and any transients respectively matched to star-like sources (defined as a source assigned a LS morphological type of “PSF”) or minor planets were removed from our search. The LS crossmatch was also used to sort the sources into three categories based on proximity to the nearest galaxy-type777We consider all LS sources with PSF types other than “PSF” (stellar) and “DUP” (extended source components) as “galaxy-type” sources. LS source: sources within 0.3” of a galaxy-type source were labeled as “A” sources, sources greater than 1.0” away from the nearest such source were labeled as “T” sources, sources falling in the middle ground between these two categories were labeled as “C” sources.
Two final cuts were then applied to limit our search to the most realistic persistent sources, requiring sources to have least two distinct observations separated by 30 minutes, and at least one high-fidelity detection (real/bogus CNN score 0.9). Of the remaining 3558 sources, 2388 of them were within 1” of a galaxy-type LS object (“A” and “C” sources); this final group composed our automated list of candidates based on our photometric DECam data. See Table LABEL:tab:vetting for a summary of cuts applied and resulting numbers of candidates at each cut.
3.2 Additional vetting
Our 2388 automatically-identified sources were then examined by eye to identify candidates for further study. Sources flagged for further investigation included those that had brightened since the time of the GW event and did not exhibit early reddening (this latter criteria was motivated by the expectation that the luminous counterpart emerges from the depths of the accretion disk, see Figure 1, and we do not expect it to redden as the optical depth decreases). Other considerations included whether there was a perceptible delay time from the GW event time (reflective of an EM counterpart needing time to escape the embedding accretion disk before being observed) and whether sources appeared similar to different known transients such as supernovae. Image stamps were also examined for each source to exclude any artifacts and the like that survived the CNN cut.
The resulting list was trimmed to remove any transients over 2 away from the GW distance posterior, as such information was available. To determine distances to each transient, the LS galaxy-type object matched to each transient was cross-matched with several galaxy catalogs (with a search radius of 1”) in search of a redshift measurement. Galaxy catalogs used included the NASA/IPAC Extragalactic Database (NED) Local Volume Sample (Cook et al., 2023) 888https://ned.ipac.caltech.edu/, the Dark Energy Spectroscopic Instrument (DESI) galaxy catalog (DESI Collaboration et al., 2023), and Quaia (Storey-Fisher et al., 2024); LS DR10.1 photometric redshifts were also included. If the host galaxy matched in multiple catalogs, then the best redshift measurement was used, preferring spectroscopic redshifts over Quaia “spectrophotometric” redshifts (Storey-Fisher et al., 2024), and the latter over photometric redshifts. Comparing the available redshift measurements to the distance posteriors (taken from the skymap HEALPix999https://healpix.sourceforge.io/ tile containing the respective transient) informed the elimination of sources beyond the 2 threshold cut, using the larger of the two uncertainties between the GW and galaxy catalog measurements. Note that transients whose host galaxy lacked any redshift measurement were not subject to this cut.
Candidates in our shortlist were photometrically classified with Parameterization of SuperNova Intrinsic Properties (ParSNIP) (Boone, 2021), assuming a model trained on the PLAsTiCC (Kessler et al., 2019) simulations; because ParSNIP uses redshift as an input parameter, this classification was limited to only those sources with a redshift measurement. When classifying a light curve, ParSNIP assigns a probability to each of a set list of transient classes so that the total probability sums to 1, with each probability reflecting the relative likeness of the light curve to each class. We note that ParSNIP does not contain a catch-all class such as “Other” for use when classifying transients of unfamiliar phenomenology: that is, if a transient is strongly dissimilar to all but one of the classes, it could receive a high classification probability for that one class, even if it not a strong match in and of itself. We consider the “TDE” ParSNIP class as the one most similar to the AGN flares of interest (all other classes concern some kind of SN, except for the KN class, whose typical timescales are considerably shorter than those we are interested in). Accordingly, in our use case we broadly interpret transients classified as “TDE” as those that are not identifiable as SN-like events, and we refer to the “TDE” class as “Non-SN” to better reflect this perspective. We find that most of our classified transients receive this classification.
The final shortlist of 23 transients surviving these cuts is shown in Table LABEL:tab:candidates. We separate several subsets from this list to identify transients that we exclude from our search via asynchronous analysis after our follow-up campaign; these subsets appear in labeled sections of the table. 6 of the transients in our shortlist showed significant brightening, but did not peak during the time they were observed (so potentially consistent with longer-term AGN variability), and so we are unable to consider their full nature with our present dataset. 9 transients, while initially interesting for further monitoring, exhibited reddening in later epochs, and so are excluded through disagreement with our assumed counterpart model that predicts a signature that becomes more blue with time. Finally, 2 transients were excluded as counterpart candidates through spectroscopic classification (see §4.2.2 and §4.2.3). The remaining 6 transients are those that cannot be excluded as counterparts to S230922g, and are listed at the top of the table. Table LABEL:tab:candidates includes host redshift information (where available), along with GW skymap localization information: the 1 distance posterior expressed as redshift under the assumed cosmology, and the 2D and 3D confidence interval (CI) as calculated with the ligo.skymap.postprocess.crossmatch.crossmatch routine101010https://lscsoft.docs.ligo.org/ligo.skymap/postprocess/crossmatch.html#ligo.skymap.postprocess.crossmatch.crossmatch (Singer et al., 2016). The transients are ranked by ascending 2D CI, s.t. the objects in the highest probability regions are listed first. The highest-probability ParSNIP class and the associated probability are listed in the last two columns of the table, where available.
AT 2023adwb | - | - | 0.136 | - | - | - | |
AT 2023uho | - | - | 0.484 | - | - | - | |
AT 2023adwp | - | - | 0.652 | - | - | - | |
AT 2023aagj | 0.184 | Specz (this work) | 0.754 | 0.829 | Non-SN | 0.944 | |
AT 2023adwt | - | - | 0.756 | - | - | - | |
AT 2023uea | Quaia SPz | 0.790 | 0.814 | Non-SN | 0.991 | ||
Transients without peak | |||||||
AT 2023adio | Quaia SPz | 0.242 | 0.363 | SNII | 0.901 | ||
AT 2023adwd | - | - | 0.285 | - | - | - | |
AT 2023adwe | LS photz | 0.285 | - | Non-SN | 0.927 | ||
AT 2023adwn | LS photz | 0.617 | 0.810 | Non-SN | 0.676 | ||
AT 2023adwq | Quaia SPz | 0.683 | 0.784 | Non-SN | 0.968 | ||
AT 2023adfo | - | - | 0.729 | - | - | - | |
Reddening transients | |||||||
AT 2023adwc | 0.216 | NED specz | 0.143 | 0.080 | Non-SN | 0.996 | |
AT 2023adwf | 0.180 | NED specz | 0.300 | 0.155 | Non-SN | 0.985 | |
AT 2023adwj | - | - | 0.429 | - | - | - | |
AT 2023uec | - | - | 0.459 | - | - | - | |
AT 2023uos | - | - | 0.475 | - | - | - | |
AT 2023adwl | - | - | 0.583 | - | - | - | |
AT 2023unj | - | - | 0.642 | - | - | - | |
AT 2023unl | 0.248 | Specz (this work) | 0.677 | 0.652 | - | - | |
AT 2023adws | 0.178 | NED specz | 0.750 | 0.827 | Non-SN | 0.928 | |
Excluded via spectra | |||||||
AT 2023aden | - | - | 0.658 | - | - | - | |
AT 2023uab | 0.128 | Specz (this work) | 0.785 | 0.554 | SNIa | 0.814 |
4 Candidate counterparts
In this section we describe candidate counterparts that received particular scrutiny during our campaign, including one we identify as the most likely to be a counterpart to S230922g. While we generally refer to candidates by their TNS name, we include the internal name for each of our transients as well, as these names were used for initial reports. Internal names are assembled from the date of discovery and RA/dec of the source; for example, the internal name of C202309242206400m275139 refers to a transient first detected on September 24, 2023 with an RA/dec of 22h 06m 40.0s, -27∘ 51m 39s (the letter between the RA and dec can be either “p” or “m”, denoting a positive or negative declination, respectively). The leading character in the object names indicates the proximity of the transient to the nearest galaxy-type LS source, in the same convention explained in §3.1 (“A” for nuclear sources, “T” for non-nuclear sources, and “C” for marginally nuclear sources). The internal names for all transients in our shortlist can be found in the subfigure titles of Figure 5.
4.1 AT 2023aagj (C2023092422006400m275139)
We consider AT 2023aagj to be the transient from our sample most likely to be an optical counterpart to S230922g. This transient is distinguished by a 1 magnitude brightening over a period of a month, remaining blue in color throughout its evolution. It is within 0.5” of the host galaxy centroid, and as such may be associated with activity in the host nucleus. Difference photometry, spectra, and sample stamps for this candidate are visible in Figure 3. ParSNIP classifies AT 2023aagj as a “TDE”/“Non-SN”-type object with a probability of 94.4%, which in this context distinguishes the transient from SNe and the like.
Two spectra of AT 2023aagj were taken before the transient faded. The initial Keck Low-Resolution Imaging Spectrograph (LRIS) spectrum, taken on December 7 (PI: Kasliwal, PID: C360), and reduced with the standard LPIPE routine (Perley, 2019), revealed ionized gas emission lines and broad line features for the H and MgII regions, indicating a possible AGN host. The redshift of the host was calculated to be 0.184, placing the transient within the 83% credible volume for the S230922g skymap and consistent at the 1.4 level with the luminosity distance posterior of the GW event. An additional spectrum of the object was taken with Gemini North Gemini Multi-Object Spectrograph (GMOS) on December 25, 2023 (PI: Cabrera, PID: GN-2023B-DD-109), using a B480 grating with 2x900s dithered exposures and a central wavelength of 6250 Å. The data were reduced with Data Reduction for Astronomy from Gemini Observatory North and South (DRAGONS) (Labrie et al., 2019), and flexure corrections were applied by re-running the reduction software without sky subtraction, and calculating a scalar offset by comparing to a sky line catalog (Rousselot et al., 2000). An interesting feature observed in the Keck spectrum is some degree of asymmetry (skewed towards the redder wavelengths) in both the H and MgII broad lines, as seen in Figure 4; however, the H asymmetry does not appear in the later Gemini North-GMOS spectrum. The latter spectrum does not cover the MgII feature, and so no comparison between the two times is possible in this regime. Given that there is no similarly-resolved archival spectrum for this source, we are not able to distinguish transient spectral features from persistent host features at this time.
We fit the Gemini spectrum using the Penalized Pixel-Fitting software (ppxf; Cappellari 2023) in order to study the possible presence of an AGN. We fit the spectrum with stellar templates from the E-MILES SSP library (Vazdekis et al., 2016) along with a number of narrow and broad Gaussian peaks corresponding to various emission and absorption lines. Using these measured gas line strengths, we place this object in the Baldwin–Phillips–Terlevich (BPT) diagram, finding that it lands in the “composite” region (according to the delineations of Kewley et al. 2001; Kauffmann et al. 2003; Kewley et al. 2006). This indicates that it is likely that the observed lines are due to a combination of star formation and AGN activity. We further measure the mass of the AGN using the fitted H velocity dispersion, which we find to be km/s. The resulting BH mass according to the method presented in (Greene & Ho, 2005) is . Note that this measurement may be inaccurate if the transient is significantly contributing to the BLR emission.
Wide-field Infrared Survey Explorer (WISE) (Wright et al., 2010a) infrared data for the host of this event is shown in Figure 3. Following Clark et al. (2024), we retrieved this data from the NASA/IPAC infrared science archive (IRSA)111111https://irsa.ipac.caltech.edu/, utilizing data from both the ALLWISE (Wright et al., 2010b) and NEOWISE Reactivation Releases (NEOWISE-R; Mainzer et al., 2011, 2014). Individual observations are combined following quality control cuts to obtain one mean magnitude per filter for each observation period, providing a 6 month cadence. Notably, there are variations in the host system of similar amplitude to our detected transient, but it must be recognized that such variations occur on much longer () timescales than the timescale of AT 2023aagj. We compare this behaviour with publicly available PanSTARRS (Flewelling et al., 2020) and DECam images at this location, which, although sparse, do not show any significant change in flux. We observe that a significant jump of mag in band 5” aperture magnitude occurred according to the SkyMapper data release (Onken et al., 2024) between MJD 57650 and 57850. It is possible that an associated optical signature existed in that time period, but was missed by the relatively sparse nature of the optical observations, and the infrared (IR) component we observed was delayed from the optical via dust reprocessing. We do not observe a rebrightening of the source in the WISE bands over the timescale of the optical transient, but there appears to be a increase in magnitude in December 2023, potentially indicating the start of a reprocessed, delayed emission from the event.
We also produce a WISE color-color plot for W1-W2 versus W2-W3, although this is not shown as W3 is only available in one epoch, which may not be representative of the entire color evolution of the object over the past years. By comparison with populations of extragalactic objects (as defined in Wright et al. 2010a) we find that while the host most clearly lies in the WISE population regions associated with spiral and luminous infrared galaxies, it is near the boundary into the Seyfert and quasar parameter spaces, and has been observed to temporarily visit these regions.
We carried out a ToO observation of AT 2023aagj with the Neil Gehrels Swift Observatory (Gehrels et al., 2004) X-ray Telescope (XRT; Burrows et al., 2005). Our observation began on 2023-12-25 at 06:02 UT with a total exposure of 3265 s in Photon Counting (PC) mode. We used the Swift-XRT automated tools (Evans et al., 2009, 2023) to analyze our data and all previous data covering the source position from May 2006 and April 2021 (totaling 7.8 ks exposure in PC mode, including the latest data). We measure a source position of RA, DEC (J2000) = , with an uncertainty radius of (90% confidence). In our latest observation the source displays a soft spectrum with all photon counts below 5 keV, similar to the previous observations. The time-averaged spectra were fit with an absorbed powerlaw model using XSPEC v12.14.0 (Arnaud, 1996) within HEASoft v6.32. The inferred photon index is relatively unconstrained but favors with hydrogen column density cm-2. The unabsorbed flux ( keV) is erg cm-2 s-1 in our latest observation. We find a factor of increase in flux between the 2006 and 2021 observations, but no significant change (a factor of ) between 2021 and 2023 (due also to the larger errors on the flux). We therefore conclude that there is no ongoing X-ray outburst from AT 2023aagj, although we note that these observations were taken months after the optical peak at a time when any additional flux from the transient may have been too faint to be detected. It is interesting to note how the WISE flare peaked around 2018, notably coincident with the significant change in X-ray flux. We speculate this may be related to a changing look AGN event, although this is hard to confirm due to the lack of an older spectrum.
4.2 Other candidates of scrutiny
We took spectra for three additional candidates beyond AT 2023aagj during our follow-up campaign. The spectrum for one transient (AT 2023unl) provides inconclusive evidence for the nature of the source; however, the source was observed to redden in later epochs, and so is disfavored by our assumed counterpart model. The spectra for the other two transients (AT 2023aden and AT 2023uab) contain SN-like features; both of these transients are subsequently excluded as possible counterparts to S230922g. The three spectra are shown in Figure 9 (Appendix A).
4.2.1 AT 2023unl (C202310042207549m253435)
The spectrum for AT 2023unl was taken by the Robert Stobie Spectrograph (RSS) (Burgh et al., 2003; Kobulnicky et al., 2003) on the South African Large Telescope (SALT) (Buckley et al., 2006). The RSSMOSPipeline (Hilton et al., 2018)121212https://github.com/mattyowl/RSSMOSPipeline was used for the reduction of SALT-RSS data. The pipeline automatically identifies and extracts one dimensional spectra using the data products delivered by the SALT observation team. Using the known redshift of the observed object and the reduced spectra, PYQSOFit (Guo et al., 2018)131313https://github.com/legolason/PyQSOFit was then used to fit and identify possible emission lines in the data. The signal-to-noise ratio (SNR) for the spectrum is relatively low, with few strong features. The present H line does not demonstrate any unusual characteristics such as asymmetry. A possible broad line feature is present around an observed wavelength of 5525 Å, albeit a chip gap appears to have ended up at the peak of the feature, and so it is uncertain how much of the observed variation is due to imperfect data reduction; that said, there is no similar feature around the second chip gap near 6575 Å. We conclude there is no conclusive evidence for this transient being in an AGN.
4.2.2 AT 2023aden (A202310262246341m291842)
A spectrum was taken for this transient using Gemini North GMOS on November 16, 2023 (PI: Cabrera, PID: GN-2023B-DD-103), and was acquired and reduced in a similar manner as the Gemini spectrum for AT 2023aagj, with a central wavelength of 7600 Å. The spectrum is blue and has strong absorption features characteristic of SNe; subsequently, we do not consider this transient as a likely counterpart to S230922g.
4.2.3 AT 2023uab (C202309242248405m134956)
We used the Double Spectrograph (DBSP) mounted at the 200-inch Hale Telescope at Palomar Observatory. For this source, we used a 1.5 arcsec slitmasks, a D55 dichroic, a blue grating of 600/4000 and red grating of 316/7500. The data were reduced using a custom PyRAF DBSP reduction pipeline (Bellm & Sesar, 2016). The spectrum exhibits clear H and H features which enable a redshift measurement of 0.128 (computed with Redrock141414https://github.com/desihub/redrock; Bailey et al., in prep.). The continuum of the spectrum peaks around 5000 Å, with several broad absorption features present around the peak. This, in concert with ParSNIP classifying the object as a type Ia SN, leads us to exclude this object as a possible counterpart.
5 Discussion
Taking our favored counterpart candidate of AT 2023aagj, we work under the assumption that the transient is related to S230922g and try to infer whether the the properties of the flare may be consistent with expectations for the counterpart signal, mechanism, and host.
What can be noted is that the persistent blue color is unlike typical supernovae, and structure function (SF) arguments as in Palmese et al. (2021) show that this transient is distinguished from typical AGN activity at the 5.9 level (corresponding to a one-sided probability ). Taking the 90% CI volume of the event as Mpc3 (as calculated with ligo.skymap) and a fiducial AGN number density of Mpc-3 (Bartos et al., 2017a; Greene & Ho, 2007), we estimate a count of 6600 AGN in the volume. Alternatively, crossmatching the Milliquas (Flesch, 2023) quasar catalog with the skymap (also with ligo.skymap) yields a 3D match of 826 objects. Given the discovery window of our follow-up campaign days (calculated as the difference in time between the second epoch and the last), we calculate the probability of a chance discovery of a similar flare as
which predicts and for the number density and Milliquas estimates, respectively.
We compare our candidate flare with the sample of Zwicky Transient Facility (ZTF) transients from Graham et al. (2023). Assuming a roughly constant spectral flux density across the - and -band frequency domain, we apply the bandpass bolometric correction of 5.15 from Duras et al. (2020) to derive a total flare energy of
where is the transmission function for the bandpass (Bessell 1990), and and are the times of the first and last observations of the flare. Fitting a Gaussian rise-exponential decay model to the -band light curve yields a rise time parameter (the standard deviation of the Gaussian) of . This places this transient in an intermediate regime between the SN and TDE populations from Graham et al. 2023 (c.f. Figure 2 in that paper) among fast TDEs and high-energy SNe. Notably, the Graham et al. 2023 proposed flare counterparts are all slower and more energetic than AT 2023aagj; this may be associated with the measurement of our flare parameters being based on difference photometry, while the ZTF light curves used in the compared study are assembled via direct photometry that includes the flux of the host AGN.
We estimate physical parameters of the system following the methodology of Graham et al. (2020), which tests the hypothesis that the flare is generated from a Bondi accreting BH as the remnant emerges from the accretion disk due to a post-merger kick. With a flare onset time delay of days from the GW trigger and total mass estimate of , under the assumption of association we can conclude one or more of the following: either (i) the kick velocity is large, or (ii) the disk is geometrically thin (aspect ratio ) or (iii) the disk is not optically thick away from a geometrically thin mid-plane, or (iv) there is a cavity in the disk due to feedback from the pre-merger BBH.
We continue by assuming the kick velocity is not specially large or small. If we assume a kick velocity (which is consistent with the peak of the prior in Varma et al. 2022) then we can constrain the approximate disk height from (Graham et al., 2023)
where is the disk height and is the mid-plane optical depth. Since , yields a factor from the square root, so . This implies a thin disk, possibly similar to the thin regions of a Thompson et al. (2005) model, which also has a lower than other models such as Sirko & Goodman (2003). The thinner regions of AGN disks are where gas damping is most efficient and consequently where mergers are more likely to occur (McKernan & Ford, 2024).
If we assume from Figure 3 that the flare begins 2 days post-S230922g (denoted by a vertical line), and the flare lasts through the spectra denoted by the blue and red lines, then the overall flare duration is days. Given , the average luminosity of the flare is , which can be parameterized as where is the Eddington luminosity. If we parameterize the Bondi accretion rate onto the merged BH () as:
where is the disk gas density and we assume the relative velocity () between the BH and the AGN disk gas is , i.e. the sound speed in the AGN disk gas is less than . We have chosen to be comparable to the typical density in a Thompson et al. (2005) model where the disk is near its thinnest ( aspect ratio ). Then , where . Interestingly, this accretion rate is within an order of magnitude of the inferred accretion rate onto embedded BBH in AGN from recent GRMHD simulations (Dittmann et al., 2024).
Associating the line asymmetry with an off-center flare implies that the signature persists in the BLR for post event, but is gone post event. A light-travel time across the full BLR of ( in rest frame), corresponds to a distance scale of , implying that the merger occurred around . This is located further out than the thinnest regions of a Sirko & Goodman (2003) model disk, so we can rule out an origin in a disk similar to that model. This location could be consistent with the thinnest () regions of a Thompson et al. (2005) model disk, which extend .
An alternative candidate for an off-center luminous flare in an AGN is a -TDE, where a star is tidally disrupted by a stellar mass BH embedded in an AGN disk. Perna et al. (2021) show that in the outer regions () of a Thompson et al. (2005) disk model, a -TDE will peak in afterglow very quickly (hours), which is inconsistent with our observations. However, if is actually a factor of a few more massive, then the diffusion time becomes longer, potentially months, which could be consistent with our observations.
We also note that the lightcurve of AT 2023aagj resembles those of afterglows in AGN disks as found in Wang et al. (2022). The rise of an afterglow from a BBH merger is not implausible, assuming that these objects are able to launch jets, as explored by Tagawa et al. (2023). Future work could investigate the possibility of fitting those afterglow models in high-density environments.
Under the assumption that AT 2023aagj is the counterpart to S230922g, our requirements are that the AGN disk must be relatively thin, similar to a Thompson et al. (2005) model, and that the accretion rate onto the kicked BH is significantly super-Eddington, but well below the Bondi rate (and requires an associated jet for the radiation to emerge). One of our most likely false positive for this event is a -TDE in the disk if our estimate of is too small by a factor of a few.
6 Conclusions
In this work we present the results of the GW-MMADS follow-up of S230922g. We discuss in detail our most likely candidate AT 2023aagj, noting it is mostly constant blue color, AGN-hosted nature, and especially the presence of asymmetry in BLR spectral features as evidence in favor of recognizing it as a counterpart to S230922g. However, we do not find the current dataset sufficient to clearly identify the transient as such.
It is worth noting that several aspects lead to an inconclusive outcome regarding the significance of this association. The variable nature of the AGN in question, showing a significant flare in the infrared and a change in X-ray flux over the past decade, exemplifies that there is some inherent activity in the host, and the discovery of a GW counterpart is only possible if a means of distinguishing the transient signal from any host variability. One possible discriminator is the observation of asymmetric components in the broad lines of the AGN spectrum, indicative of an off-center phenomenon; however, other sources such as disk winds and tilted dust obscuration may cause similar profiles. A smoking gun for the BBH association would be the evolution of the asymmetric component in sync with the transient light curve, with the asymmetry shifting towards bluer wavelengths and eventually disappearing as the transient fades. For AT 2023aagj, due to the lack of wavelength coverage for the Gemini spectrum and subsequent lack of multiple observations of MgII, a chip gap on H in the Keck spectrum, the telluric region blueward of H, and the target setting in December, the extraction of such an evolution is not feasible with the present data. A spectrum of this object in 2024 as the target rises again may be informative about the nature of the asymmetry. A persistent line asymmetry long after the merger event may point to similarly persistent asymmetric illumination in this source, such as a warped disk, and could rule out the association between AT 2023aagj and S230922g. Finally, given the blue color and the ParSNIP prediction, we cannot exclude that this transient may have originated from a TDE.
[…]
Acknowledgements
TC, AP, and LH acknowledge that this material is based upon work supported by NSF Grant No. 2308193. BO gratefully acknowledges support from the McWilliams Postdoctoral Fellowship at Carnegie Mellon University. BM & KESF are supported by NSF AST-2206096 and NSF AST-1831415 and Simons Foundation Grant 533845 as well as Simons Foundation sabbatical support. The Flatiron Institute is supported by the Simons Foundation. AP thanks Rosalba Perna, Armin Rest, and Stephen Smartt for useful discussion.
This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award HEP-ERCAP0029208 and HEP-ERCAP0022871. This work used resources on the Vera Cluster at the Pittsburgh Supercomputing Center (PSC). We thank T.J. Olesky and the PSC staff for help with setting up our software on the Vera Cluster. MWC acknowledges support from the National Science Foundation with grant numbers PHY-2308862 and PHY-2117997.
This project used data obtained with the Dark Energy Camera (DECam), which was constructed by the Dark Energy Survey (DES) collaboration. Funding for the DES Projects has been provided by the US Department of Energy, the US National Science Foundation, the Ministry of Science and Education of Spain, the Science and Technology Facilities Council of the United Kingdom, the Higher Education Funding Council for England, the National Center for Supercomputing Applications at the University of Illinois at Urbana-Champaign, the Kavli Institute for Cosmological Physics at the University of Chicago, Center for Cosmology and Astro-Particle Physics at the Ohio State University, the Mitchell Institute for Fundamental Physics and Astronomy at Texas A&M University, Financiadora de Estudos e Projetos, Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro, Conselho Nacional de Desenvolvimento Científico e Tecnológico and the Ministério da Ciência, Tecnologia e Inovação, the Deutsche Forschungsgemeinschaft and the Collaborating Institutions in the Dark Energy Survey.
The Collaborating Institutions are Argonne National Laboratory, the University of California at Santa Cruz, the University of Cambridge, Centro de Investigaciones Enèrgeticas, Medioambientales y Tecnològicas–Madrid, the University of Chicago, University College London, the DES-Brazil Consortium, the University of Edinburgh, the Eidgenössische Technische Hochschule (ETH) Zürich, Fermi National Accelerator Laboratory, the University of Illinois at Urbana-Champaign, the Institut de Ciéncies de l’Espai (IEEC/CSIC), the Institut de Física d’Altes Energies, Lawrence Berkeley National Laboratory, the Ludwig-Maximilians Universit at M unchen and the associated Excellence Cluster Universe, the University of Michigan, NSF’s NOIRLab, the University of Nottingham, the Ohio State University, the OzDES Membership Consortium, the University of Pennsylvania, the University of Portsmouth, SLAC National Accelerator Laboratory, Stanford University, the University of Sussex, and Texas A&M University.
Based on observations at Cerro Tololo Inter-American Observatory, NSF’s NOIRLab (NOIRLab Prop. ID 2022B-715089; PI: Palmese; 2023B-851374 , PI: Andreoni & Palmese; 2023B-735801, PI: Palmese & Wang), which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. We thank Kathy Vivas, Alfredo Zenteno, and CTIO staff for their support with DECam observations.
Based on observations obtained at the international Gemini Observatory (Prop. ID GN-2023B-DD-103, PI: Cabrera; GN-2023B-109, PI: Cabrera), a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation on behalf of the Gemini Observatory partnership: the U.S. National Science Foundation (United States), National Research Council (Canada), Agencia Nacional de Investigación y Desarrollo (Chile), Ministerio de Ciencia, Tecnología e Innovación (Argentina), Ministério da Ciência, Tecnologia, Inovações e Comunicações (Brazil), and Korea Astronomy and Space Science Institute (Republic of Korea). Gemini data were processed using DRAGONS (Data Reduction for Astronomy from Gemini Observatory North and South).
This work was enabled by observations made from the Gemini North telescope, located within the Maunakea Science Reserve and adjacent to the summit of Maunakea. We are grateful for the privilege of observing the Universe from a place that is unique in both its astronomical quality and its cultural significance.
Some of the observations reported in this paper were obtained with the Southern African Large Telescope (SALT).
The Legacy Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS; Proposal ID 2014B-0404; PIs: David Schlegel and Arjun Dey), the Beijing-Arizona Sky Survey (BASS; NOAO Prop. ID 2015A-0801; PIs: Zhou Xu and Xiaohui Fan), and the Mayall z-band Legacy Survey (MzLS; Prop. ID 2016A-0453; PI: Arjun Dey). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. Pipeline processing and analyses of the data were supported by NOIRLab and the Lawrence Berkeley National Laboratory (LBNL). The Legacy Surveys project is honored to be permitted to conduct astronomical research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. LBNL is managed by the Regents of the University of California under contract to the U.S. Department of Energy.
BASS is a key project of the Telescope Access Program (TAP), which has been funded by the National Astronomical Observatories of China, the Chinese Academy of Sciences (the Strategic Priority Research Program “The Emergence of Cosmological Structures” Grant XDB09000000), and the Special Fund for Astronomy from the Ministry of Finance. The BASS is also supported by the External Cooperation Program of Chinese Academy of Sciences (Grant 114A11KYSB20160057), and Chinese National Natural Science Foundation (Grant 12120101003, 11433005).
The Legacy Survey team makes use of data products from the Near-Earth Object Wide-field Infrared Survey Explorer (NEOWISE), which is a project of the Jet Propulsion Laboratory/California Institute of Technology. NEOWISE is funded by the National Aeronautics and Space Administration.
The Legacy Surveys imaging of the DESI footprint is supported by the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under Contract No. DE-AC02-05CH1123, by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract; and by the U.S. National Science Foundation, Division of Astronomical Sciences under Contract No. AST-0950945 to NOAO.
This research has made use of the NASA/IPAC Extragalactic Database (NED), which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology.
Appendix A Additional data
This section contains light curves (Figures 5-8) and spectra (Figure 9) for all candidates other than AT 2023aagj.
For some transients we took additional data with the Three Channel Camera (3KK) (Lang-Bardl et al., 2016) at the 2.1m-Fraunhofer Wendelstein Telescope (Hopp et al., 2014) located in the German alps. The instrument is capable of taking three images at the same time in two optical passbands and one near-infrareds (NIRs) band. For the monitoring of the transients we used the , and bands. The data were reduced with a pipeline developed at the University Observatory Munich. For the detrending of the images the pipeline makes use of the tools from Gössl & Riffeser (2002). Then the images are calibrated and coadded using SExtractor (Bertin & Arnouts, 1996), SCAMP (Bertin, 2006) and SWarp (Bertin et al., 2002), difference photometry is conducted with SFFT (Hu et al., 2022), and extinction corrections are applied as described in §3.1. During this process, the astrometric solution is computed against the Gaia EDR3 catalog (Gaia Collaboration et al., 2016, 2021) and the zero points for the optical bands are calibrated with the PanSTARRS 1s (PS1s) catalog (Magnier et al., 2013). The NIRs zero points are matched to the Two Micron All Sky Surveys (2MASSs) catalog (Skrutskie et al., 2006) and converted to AB magnitudes (Blanton & Roweis, 2007). Wendelstein data are plotted as outlined diamonds in Figure 5, with , , and band data plotted in green, yellow, and red, respectively.